On a coordinate grid, your campsite is located at (2,-10) and the next checkpoint station is located at (0,-3). Each unit in the coordinate plane represents 1.4 miles. Find the distance you have to hike to reach the next checkpoint station.
1
Expert's answer
2020-10-07T15:52:17-0400

Location of Campsite = (2, -10)

Location of Checkpoint Station = (0, -3)

."\begin {aligned}
\therefore
\end {aligned}" x1 = 2 ; y1 = -10

x2 = 0 ; y2 = -3


1 unit = 1.4miles


let, the distance between campsite and checkpoint station = r


from, distance between 2 points = "\begin {aligned}
\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}
\end {aligned}"

r = "\begin {aligned}
\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}
\end {aligned}"

r = "\begin {aligned}
\sqrt{(0-2)^2+(-3-(-10))^2}
\end {aligned}"

r = "\begin {aligned}
\sqrt{(-2)^2+(-3+10)^2}
\end {aligned}"

r = "\begin {aligned}
\sqrt{4+(-7)^2}
\end {aligned}"

r = "\begin {aligned}
\sqrt{4+49}
\end {aligned}"

r = "\begin {aligned}
\sqrt{53}
\end {aligned}"

r = 7.28 units

"\begin {aligned}
\therefore
\end {aligned}" the distance between campsite and checkpoint station = 7.28 units


However since 1 unit = 1.4 miles,

7.28 units = 1.4 "\begin {aligned}
\times
\end {aligned}" 7.28 = 10.192 miles


"\begin {aligned}
\therefore
\end {aligned}" the distance between campsite and checkpoint station = 10.192 miles